An object is thrown straight up with an initial velocity of 80m/s. What is its maximum height that it reaches?

a. 32.62 m b.80 m c. 99.38 m d. 326.20 m




1
Expert's answer
2021-04-23T11:08:14-0400
"h=\frac{v^2}{2g}\\h=\frac{80^2}{2(9.8)}=326.20\ m"

Answer: d. 326.20 m